# Communication-Aware Dissipative Control for Networks of Heterogeneous Nonlinear Agents

### v1

@@ 4.3 Initialization
Initial feasible points of Equation 9 or 12, denoted by $\mathbf{P}^{0}$, $\mathbf{K}^{0}$, $\mathbf{Q}^{0}$, $\mathbf{S}^{0}$, $\mathbf{R}^{0}$, $\widehat{\mathbf{Q}}^{0}$, $\widehat{\mathbf{S}}^{0}$, $\widehat{\mathbf{R}}^{0}$, $\mathbf{G}^{0}$, and $\mathbf{F}^{0}$, are required to solve Equation 12. Heuristically, localized controllers $\mathbf{K}_{i}\in\mathbb{R}$ for $i\in\mathbb{N}_{N}$ can be initialized based on classical control methods, such as PD control, and a feasible point of Equation 9 can then be obtained with these fixed local controllers. If the constraints are not satisfied for a chosen localized controller, the feasibility check is repeated with increased controller gains until a feasible point is found. Once a feasible set of local controller is obtained, solving Equation 12 results in the initial feasible point corresponding to a centralized initial controller. Alternatively, the iterative relaxation strategy proposed in (Warner and Scruggs, 2017) can also be employed to obtain an initial feasible solution. For $\mathbf{G}^{0}$ and $\mathbf{F}^{0}$, any matrices satisfying conditions in Corollary 1 can be used, and $\mathbf{I}$ is commonly used for the initialization.
@@ 4.4 Sparsity Promotion
where $W_{ij}$ are non-negative weight. If $W_{ij}$ are inversely proportional to $\left\|\mathbf{K}_{ij}^{0}{+}\delta{\mathbf{K}}_{ij}\right\|_{F}$, the weighted $\ell_{1}$ norm becomes equivalent to the cardinality penalty. However, this exact formulation cannot be implemented, since $W_{ij}$ depend on the unknown variables (Lin et al., 2013). Instead, the structure of the initial feasible controller $\mathbf{K}^{0}$ can be used to define $W_{ij}$ as $\left\|\mathbf{K}^{0}_{ij}\right\|_{F}^{-1}$ when $\left\|\mathbf{K}^{0}_{ij}\right\|_{F}{\neq}0$ or $\epsilon_{l}^{-1}$ otherwise. When this weighted $\ell_{1}$ norm is used, the penalty function becomes convex, allowing Equation 13 to be solved directly by standard optimization solvers without relying on specialized algorithms.
Lin et al. (2013) proposed solving the problem with cardinality penalty using the ADMM algorithm. Although ADMM does not guarantee convergence for nonconvex objectives, extensive computational experience has shown that it performs well in practice when the augmented Lagrangian parameter is sufficiently large (Boyd et al., 2011).
The same strategy in Lin et al. (2013) is applied to Equation 13 using cardinality penalty function
@@
By introducing the indicator function $I_{\mathbf{N}}$ and an auxiliary variable $\mathbf{Z}$, where $I_{\mathbf{N}}$ takes 0 if $\mathbf{N}\preceq 0$ and $\infty$ otherwise, Equation 13 is equivalently written as
@@
where $J(\delta{\mathbf{P}}){=}\text{Trace}(\mathbf{P}^{0}{+}\delta{\mathbf{P}})$. The goal of ADMM is to solve
@@
where $\bm{\Lambda}$ is the dual variable, $k$ is the iteration index, $\rho>0$ is the augmented Lagrangian parameter. If Equation 15 is used as the penalty, the block component of the unique solution of Equation 16b is
$\mathbf{Z}_{ij}^{k+1}{=}\begin{cases}\mathbf{V}_{ij},&\left\|\mathbf{V}_{ij}\right\|_{F}{>}\sqrt{\frac{2\gamma}{\rho}}\\ \mathbf{0},&o.w.\end{cases}$
where $\mathbf{V}{=}\mathbf{K}^{0}{+}\delta{\mathbf{K}}^{k+1}{+}\bm{\Lambda}^{k}$ (Lin et al., 2013). The stopping criteria of ADMM are
$r_{p}{=}\frac{\left\|\mathbf{K}^{0}{+}\delta{\mathbf{K}}^{k}{-}\mathbf{Z}^{k}\right\|_{F}}{\left\|\mathbf{Z}^{k}\right\|_{F}}{\leq}\epsilon_{p}\text{ and }r_{d}{=}\frac{\left\|\mathbf{Z}^{k}{-}\mathbf{Z}^{k-1}\right\|_{F}}{\left\|\mathbf{Z}^{k}\right\|_{F}}{\leq}\epsilon_{d}.$
After solving Equation 13, the controller is updated to $\mathbf{K}_{1}=\mathbf{K}^{0}+\delta{\mathbf{K}}^{\star}$, where $\delta{\mathbf{K}}^{\star}$ is the optimal perturbation obtained from the solution. The matrices $\mathbf{Q}^{1}$, $\mathbf{S}^{1}$, $\mathbf{R}^{1}$, $\widehat{\mathbf{Q}}^{1}$, $\widehat{\mathbf{S}}^{1}$, $\widehat{\mathbf{R}}^{1}$, $\mathbf{Y}^{1}$, and $\mathbf{G}^{1}$ are updated accordingly, and $\mathbf{F}^{1}=\mathbf{F}^{\star}$. The subspace $\mathcal{K}$ is then defined as the set of block matrices that share the same sparsity pattern as $\mathbf{K}^{1}$.
@@ 4.5 Structured ICO
The process described in Section 4.4 results in the solution of Equation 13, which represents a linear convex relaxation of Equation 9 obtained by linearizing around the initial feasible points. This implies that the resulting solutions are not necessarily local optima of Equation 9. Intuitively, the closed-loop performance can be further improved by iteratively solving Equation 13 by using the feasible points obtained from the previous iteration as the initial points. However, such an approach does not preserve the convergence guarantees of ICO, since the cardinality operator is pointwise discontinuous at zero, and no approximation of the cardinality can simultaneously be an overbound, convex, and nonconservative at zero (LoCicero and Bridgeman, 2022). Therefore, instead of iteratively solving Equation 13, we consider
@@
3:if $g(\mathbf{K}^{0}+\delta{\mathbf{K}})$ follows Equation 14 then
4:  Find $\delta{\mathbf{K}}^{\star}$ by solving Equation 13
5:else if $g(\mathbf{K}^{0}+\delta{\mathbf{K}})$ follows Equation 15 then
@@
7:   Find $\delta{\mathbf{K}}^{k+1}$ by solving Equation 16a
8:   Find $\mathbf{Z}^{k+1}$ by solving Equation 16b
9:   Find $\bm{\Lambda}^{k+1}$ by solving Equation 16c
@@
15:  Solve Equation 17 using $\mathbf{K}^{k}$, $\mathbf{Y}^{k}$, $\mathbf{Q}^{k}$, $\mathbf{S}^{k}$, $\mathbf{R}^{k}$, $\widehat{\mathbf{Q}}^{k}$, $\widehat{\mathbf{S}}^{k}$, $\widehat{\mathbf{R}}^{k}$, $\mathbf{G}^{k}$, and $\mathbf{F}^{k}$
@@ 4.6 Convergence of Algorithm 1
Algorithm 1 consists of two major stages, sparsity promotion, described in Section 4.4, and ICO, described in Section 4.5. The ICO is always feasible and converges to a local optimum, since it retains at least one feasible point, which is the solution of the previous iteration, and the cost of the problem in each iteration is guaranteed to be no greater than that of this feasible point. In contrast, the convergence behavior of the sparsity-promotion stage depends on the choice of the penalty function. When the weighted $\ell_{1}$ norm is employed, the algorithm converges to the a specific structured controller because the problem is formulated as a SDP, which can be solved using standard convex optimization solvers. On the other hand, when the cardinality function is used, there is no convergence guarantee due to its inherent nonconvexity. Nevertheless, extensive computational studies have shown that the ADMM algorithm performs well in practice when the value of $\rho$ is sufficiently large (Boyd et al., 2011).
@@ 5 Numerical Examples
The actual dynamics $\mathbf{A}_{ii}$ of each agent are uniformly distributed within $\pm 30\%$ of their nominal values $\widehat{\mathbf{A}}_{ii}$. Accordingly, each agent’s dynamics can be modeled as an LTI system with polytopic uncertainty, represented by a polyhedron with 16 vertices, where each vertex corresponds to the case in which one parameter of $\mathbf{A}_{ii}$ attains either its maximum or minimum value within the uniform distribution. Since some agents are unstable without control, (Walsh and Forbes, 2019, Equation 18) cannot be applied because it requires stability of the open-loop systems. Therefore, Lemma 2 is used, with $\mathbf{A}^{j}$ denoting the 8 vertices of the polytope and $\mathbf{B}^{j}=\mathbf{I}_{2}$ for all $j\in\mathbb{N}_{8}$.
@@
The operator $\mathscr{G}:\mathcal{L}_{2e}^{m}\mapsto\mathcal{L}_{2e}^{l}$ defined by $\dot{\textbf{x}}(t)=\mathbf{A}(t)\mathbf{x}(t)+\mathbf{B}(t)\mathbf{u}(t)$ and $\mathbf{y}(t)=\mathbf{x}(t)$, where $\mathbf{A}(t)=\sum_{j=1}^{v}\zeta_{j}(t)\mathbf{A}^{j}$ and $\mathbf{B}(t)=\sum_{j=1}^{v}\zeta_{j}(t)\mathbf{B}^{j}$ with $\sum_{j=1}^{v}\zeta_{j}(t)=1$ for all $t\in\mathbb{R}_{+}$, is $QSR$-dissipative if there exist a matrix $\mathbf{P}\succ 0$ and matrices $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}$ such that
$\begin{bmatrix}\text{He}(\mathbf{P}\mathbf{A}^{j})-\mathbf{Q}&\mathbf{P}\mathbf{B}^{j}-\mathbf{S}\\ (\mathbf{P}\mathbf{B}^{j})^{T}-\mathbf{S}^{T}&-\mathbf{R}\end{bmatrix}\preceq 0,\quad\forall j\in\mathbb{N}_{v}$ (18)
@@
The proof proceeds by verifying Lemma 1. The operator is $QSR$ dissipative if Lemma 1 holds for $(\mathbf{A}(t),\mathbf{B}(t),\mathbf{I},\mathbf{0})$. When $\mathbf{C}=\mathbf{I}$ and $\mathbf{D}=\mathbf{0}$, Equation 2 becomes affine in $\mathbf{A}(t)$ and $\mathbf{B}(t)$. Therefore,
$\sum_{j=1}^{v}\zeta(t)\begin{bmatrix}\text{He}(\mathbf{P}\mathbf{A}^{j})-\mathbf{Q}&\mathbf{P}\mathbf{B}^{j}-\mathbf{S}\\ (\mathbf{P}\mathbf{B}^{j})^{T}-\mathbf{S}^{T}&-\mathbf{R}\end{bmatrix}\preceq 0$
@@
The initial local controller for all agents is set to [1000 100][1000\;100]. To apply Algorithm 1, the parameters are chosen as $\mathbf{Q}_{lq}{=}100\mathbf{I}_{20}$, $\mathbf{R}_{lq}{=}\mathbf{I}_{20}$, $\epsilon_{l}{=}10^{-3}$, $\epsilon_{p}{=}\epsilon_{d}{=}10^{-4}$, $\rho{=}100$, and $\epsilon{=}10^{-3}$. The weighting factor $\gamma$ is varied from 100 to 300 in increments of 5 for the weighted $\ell_{1}$ norm penalty, and from 335 to 1270 in increments of 5 for the cardinality penalty. For the comparison, the approaches in (Lin et al., 2013; Lian et al., 2018) are implemented for the same system dynamics. The same procedure as in (Lian et al., 2018, Chap. 5) is used to compute the uncertainty input and output channels, with the $\mathcal{H}_{\infty}$ constraint parameter $\gamma$ set to 1.25. After obtaining the controllers, stability is evaluated over 200 randomly generated true dynamics of each agent, as well as over 256 vertex combinations constructed from the 16 vertex matrices associated with the stable and unstable agents. In the latter case, the stability analysis is performed under the simplifying assumption that all stable agents follow the same vertex matrix from the polytope corresponding to stable agents, and all unstable agents follow the same vertex matrix from the polyhedron corresponding to unstable agents.
@@
The results in Figure 2, 3 and 4 show the LQ performance, stability, and sparse structure of the network controllers resulting from the proposed method and reference approaches in (Lin et al., 2013; Lian et al., 2018). In all cases, the sparsity-promotion method converges to a sparse controller structure. The approach in (Lin et al., 2013) achieves the best LQ performance but can lead to instability due to the uncertain nature of each agent. Both versions of the proposed method, as well as the approach in (Lian et al., 2018), maintain stability under uncertain dynamics and for all vertex combinations of the polytopic uncertainty, with the proposed methods exhibiting superior LQ performance.
@@
The total computation time is also evaluated for a larger network with $N=50$ as in Table 1. In this case, $\mathcal{H}_{\infty}$ constraint parameter $\gamma$ is set to $8$. Among the stabilizable algorithms, the proposed method with weighted $\ell_{1}$ norm yields the shortest computation time, as it does not require any iterative procedure to determine a sparse communication topology.

### v2

@@ 4.3 Initialization
Initial feasible points of ()9 or ()12, denoted by $\mathbf{P}^{0}$, $\mathbf{K}^{0}$, $\mathbf{Q}^{0}$, $\mathbf{S}^{0}$, $\mathbf{R}^{0}$, $\widehat{\mathbf{Q}}^{0}$, $\widehat{\mathbf{S}}^{0}$, $\widehat{\mathbf{R}}^{0}$, $\mathbf{G}^{0}$, and $\mathbf{F}^{0}$, are required to solve ()12. Heuristically, localized controllers $\mathbf{K}_{i}{\in}\mathbb{R}$ for $i{\in}\mathbb{N}_{N}$ can be initialized based on classical control methods, such as PD control, and a feasible point of ()9 can then be obtained with these fixed local controllers. If the plant agents are stable, the strategy in (LoCicero and Bridgeman, 2025, Section 6) can be applied. If the constraints are not satisfied for a chosen localized controller, the feasibility check is repeated with increased controller gains until a feasible point is found. Once a feasible set of local controller is obtained, solving ()12 results in the initial feasible point corresponding to a centralized initial controller. Alternatively, the iterative relaxation strategy proposed in (Warner and Scruggs, 2017) can also be employed to obtain an initial feasible solution. For $\mathbf{G}^{0}$ and $\mathbf{F}^{0}$, any matrices satisfying conditions in Corollary 1 can be used, and $\mathbf{I}$ is commonly used for the initialization.
@@ 4.4 Sparsity Promotion
where $W_{ij}$ are non-negative weights. If $W_{ij}$ are inversely proportional to $\left\|\mathbf{K}_{ij}^{0}{+}\delta{\mathbf{K}}_{ij}\right\|_{F}$, the weighted $\ell_{1}$ norm becomes equivalent to the cardinality penalty. However, $W_{ij}$ cannot depend on the unknown variables (Lin et al., 2013). Instead, the structure of the initial feasible controller $\mathbf{K}^{0}$ can be used to define $W_{ij}{=}\min\{\left\|\mathbf{K}^{0}_{ij}\right\|_{F}^{-1}{,}\ \epsilon_{l}^{-1}\}$. When this weighted $\ell_{1}$ norm is used, the penalty function becomes convex, so ()13 can be solved numerically.
Lin et al. (2013) proposed to solve the problem with a cardinality penalty using ADMM. Although ADMM does not guarantee convergence for nonconvex objectives, extensive computational experience has shown that it performs well in practice when the augmented Lagrangian parameter is sufficiently large (Boyd et al., 2011). We therefore adopt the strategy of Lin et al. (2013) for ()13, using
@@
By introducing the indicator function $I_{\mathbf{N}}$ and an auxiliary variable $\mathbf{Z}$, where $I_{\mathbf{N}}$ takes 0 if $\mathbf{N}\preceq 0$ and $\infty$ otherwise, ()13 is equivalently written as
@@
where $J(\delta{\mathbf{P}}){=}\text{Trace}(\mathbf{P}^{0}{+}\delta{\mathbf{P}})$. ADMM then solves
@@
where $\bm{\Lambda}$ is the dual variable, $k$ is the iteration index, and $\rho{>}0$ is the augmented Lagrangian parameter. If ()15 is used as the penalty, the block component of the unique solution to ()16b, $\mathbf{Z}_{ij}^{k+1}$, is $\mathbf{V}_{ij}$ if $\left\|\mathbf{V}_{ij}\right\|_{F}{>}\sqrt{2\gamma/\rho}$ and $\mathbf{0}$ otherwise, where $\mathbf{V}{=}\mathbf{K}^{0}{+}\delta{\mathbf{K}}^{k+1}{+}\bm{\Lambda}^{k}$ (Lin et al., 2013). The stopping criteria of ADMM are $r_{p}{=}\frac{\left\|\mathbf{K}^{0}{+}\delta{\mathbf{K}}^{k}{-}\mathbf{Z}^{k}\right\|_{F}}{\left\|\mathbf{Z}^{k}\right\|_{F}}{\leq}\epsilon_{p}$ and $r_{d}{=}\frac{\left\|\mathbf{Z}^{k}{-}\mathbf{Z}^{k-1}\right\|_{F}}{\left\|\mathbf{Z}^{k}\right\|_{F}}{\leq}\epsilon_{d}$.
After solving ()13, the controller is updated to $\mathbf{K}_{1}{=}\mathbf{K}^{0}+\delta{\mathbf{K}}^{\star}$, where $\delta{\mathbf{K}}^{\star}$ is the optimal perturbation obtained from the solution. The matrices $\mathbf{Q}^{1}$, $\mathbf{S}^{1}$, $\mathbf{R}^{1}$, $\widehat{\mathbf{Q}}^{1}$, $\widehat{\mathbf{S}}^{1}$, $\widehat{\mathbf{R}}^{1}$, $\mathbf{Y}^{1}$, and $\mathbf{G}^{1}$ are updated accordingly, and $\mathbf{F}^{1}{=}\mathbf{F}^{\star}$. The subspace $\mathcal{K}$ is then defined as the set of block matrices that share the same sparsity pattern as $\mathbf{K}^{1}$.
@@ 4.5 Structured ICO
The process described in Section 4.4 solves ()13, which represents a conservative convexification of ()9 about the initial feasible points. This implies that the resulting solutions are not necessarily local optima of ()9. Intuitively, the closed-loop performance can be further improved by convexifying about the new solution points and re-solving ()13. However, such an approach does not preserve the convergence guarantees of ICO, since the cardinality operator is pointwise discontinuous at zero, and no approximation of the cardinality can simultaneously be an overbound, convex, and nonconservative at zero (LoCicero and Bridgeman, 2022). Therefore, instead of iteratively solving ()13, we consider
@@
3:if $g(\mathbf{K}^{0}+\delta{\mathbf{K}})$ follows ()14 then
4:  Find $\delta{\mathbf{K}}^{\star}$ by solving ()13
5:else if $g(\mathbf{K}^{0}+\delta{\mathbf{K}})$ follows ()15 then
@@
7:   Find $\delta{\mathbf{K}}^{k+1}$ by solving ()16a
8:   Find $\mathbf{Z}^{k+1}$ by solving ()16b
9:   Find $\bm{\Lambda}^{k+1}$ by solving ()16c
@@
15:  Solve ()17 using $\mathbf{K}^{k}$, $\mathbf{Y}^{k}$, $\mathbf{Q}^{k}$, $\mathbf{S}^{k}$, $\mathbf{R}^{k}$, $\widehat{\mathbf{Q}}^{k}$, $\widehat{\mathbf{S}}^{k}$, $\widehat{\mathbf{R}}^{k}$, $\mathbf{G}^{k}$, and $\mathbf{F}^{k}$
@@ 4.6 Convergence of Algorithm 1
Algorithm 1 combines two major stages, sparsity promotion, described in Section 4.4, and ICO, described in Section 4.5. The ICO is always feasible and converges to a local optimum, since it retains at least one feasible point, which is the solution of the previous iteration, and the cost of the problem in each iteration is guaranteed to be no greater than that of this feasible point. In contrast, the convergence behavior of the sparsity-promotion stage depends on the choice of the penalty function. When the weighted $\ell_{1}$ norm is employed, the algorithm converges to the a specific structured controller because the problem is formulated as a SDP, which can be solved using standard convex optimization solvers. On the other hand, when the cardinality function is used, there is no convergence guarantee due to its inherent nonconvexity.
@@ 5 Numerical Examples
The actual dynamics $\mathbf{A}_{ii}$ of each agent are uniformly distributed within $\pm 30\%$ of their nominal values $\widehat{\mathbf{A}}_{ii}$. Accordingly, each agent’s dynamics can be modeled as an LTI system with polytopic uncertainty, represented by a polyhedron with 16 vertices, where each vertex corresponds to the case in which one parameter of $\mathbf{A}_{ii}$ attains either its maximum or minimum value within the uniform distribution. Since some agents are unstable without control, (Walsh and Forbes, 2019, Equation 18) cannot be applied because it requires open-loop stability. Therefore, Lemma 2 is used, with $\mathbf{A}^{j}$ denoting the 8 vertices of the polytope and $\mathbf{B}^{j}=\mathbf{I}_{2}$ for all $j\in\mathbb{N}_{8}$.
@@
The operator $\mathscr{G}{:}\mathcal{L}_{2e}^{m}{\mapsto}\mathcal{L}_{2e}^{l}$ defined by $\dot{\textbf{x}}(t)=\mathbf{A}\mathbf{x}(t){+}\mathbf{B}\mathbf{u}(t)$ and $\mathbf{y}(t){=}\mathbf{x}(t)$, where $\mathbf{A}{=}\sum_{j=1}^{v}\zeta_{j}\mathbf{A}^{j}$ and $\mathbf{B}{=}\sum_{j=1}^{v}\hskip-2.0pt\zeta_{j}\mathbf{B}^{j}$ with $\sum_{j=1}^{v}\hskip-2.0pt\zeta_{j}{=}1$, is $QSR$-dissipative if there exist a matrix $\mathbf{P}{\succ}0$ and matrices $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}$ such that
$\begin{bmatrix}\text{He}(\mathbf{P}\mathbf{A}^{j})-\mathbf{Q}&\mathbf{P}\mathbf{B}^{j}-\mathbf{S}\\ (\mathbf{P}\mathbf{B}^{j})^{T}-\mathbf{S}^{T}&-\mathbf{R}\end{bmatrix}\preceq 0,\quad\forall j\in\mathbb{N}_{v}.$ (18)
@@
The proof proceeds by verifying Lemma 1. The operator is $QSR$ dissipative if Lemma 1 holds for $(\mathbf{A},\mathbf{B},\mathbf{I},\mathbf{0})$. Then ()2 is affine in $\mathbf{A}$ and $\mathbf{B}$, becoming
$\sum_{j=1}^{v}\zeta(t)\begin{bmatrix}\text{He}(\mathbf{P}\mathbf{A}^{j})-\mathbf{Q}&\mathbf{P}\mathbf{B}^{j}-\mathbf{S}\\ (\mathbf{P}\mathbf{B}^{j})^{T}-\mathbf{S}^{T}&-\mathbf{R}\end{bmatrix}\preceq 0.$
@@
The initial local controller gain for all agents is set to [1000 100][1000\;100]. To apply Algorithm 1, the parameters are chosen as $\mathbf{Q}_{lq}{=}100\mathbf{I}_{20}$, $\mathbf{R}_{lq}{=}\mathbf{I}_{20}$, $\epsilon_{l}{=}10^{-3}$, $\epsilon_{p}{=}\epsilon_{d}{=}10^{-4}$, $\rho{=}100$, and $\epsilon{=}10^{-3}$. The weighting factor $\gamma$ is varied from 100 to 300 in increments of 5 for the weighted $\ell_{1}$ norm penalty, and from 335 to 1270 in increments of 5 for the cardinality penalty. For the comparison, the approaches in (Lin et al., 2013; Lian et al., 2018) are implemented for the same system dynamics. The same procedure as in (Lian et al., 2018, Chap. 5) is used to compute the uncertainty input and output channels, with the $\mathcal{H}_{\infty}$ constraint parameter $\gamma$ set to 1.25. After obtaining the controllers, stability is evaluated over 200 randomly generated true dynamics of each agent, as well as over 256 vertex combinations constructed from the 16 vertex matrices associated with the stable and unstable agents. In the latter case, the stability analysis is performed under the simplifying assumption that all stable agents follow the same vertex matrix from the polytope corresponding to stable agents, and all unstable agents follow the same vertex matrix from the polyhedron corresponding to unstable agents.
@@
The results in Figure 2, Figure 3, and Figure 4 show the LQ performance, stability, and sparse structure of the network controllers resulting from the proposed method and reference approaches in (Lin et al., 2013; Lian et al., 2018). In all cases, the sparsity-promotion method converges to a sparse controller structure. The approach in (Lin et al., 2013) achieves the best LQ performance, but can lead to instability due to the agent uncertainty it neglects. Both versions of the proposed method, as well as the approach in (Lian et al., 2018), maintain stability under uncertain dynamics and for all vertex combinations of the polytopic uncertainty, with the proposed methods exhibiting superior LQ performance.
@@
The total computation time is also evaluated for a larger network with N=50N{=}50 as in Table 1. In this case, $\mathcal{H}_{\infty}$ constraint parameter $\gamma$ is set to $8$. Among the stabilizable algorithms, the proposed method with weighted $\ell_{1}$ norm yields the shortest computation time, as it does not require any iterative procedure to determine a sparse communication topology.